OpenStudy (aroub):

evaluate lim x-> pi/3 (tanx/x) Answer=1?

OpenStudy (anonymous):

Really? How you got that??

OpenStudy (anonymous):

tan(x) at x = pi/3 is a finite value.. x = pi/3 is also finite.. So, you can directly plug in the value of x there..

OpenStudy (anonymous):

\[\tan(\frac{\pi}{3}) = ??\]

OpenStudy (aroub):

\[\sqrt{3}\]

OpenStudy (aroub):

Thats the answer?

OpenStudy (anonymous):

No...

OpenStudy (anonymous):

you have tan(x) divided by x too.. So divide tan(pi/3) with pi/3

OpenStudy (anonymous):

\[\lim_{x \rightarrow \frac{\pi}{3}} (\frac{\tan(x)}{3}) \implies \frac{\tan(\frac{\pi}{3})}{\frac{\pi}{3}}\]

OpenStudy (aroub):

Well I did, \[\frac{ \frac{ sinx }{ cosx} }{ x } = \frac{ sinx }{ x }\left[ \frac{ 1 }{ cosx} \right] =1\]

OpenStudy (anonymous):

\[\lim_{x \rightarrow \frac{\pi}{3}} (\frac{\tan(x)}{x}) \implies \frac{\tan(\frac{\pi}{3})}{\frac{\pi}{3}}\]

OpenStudy (aroub):

I dont know if what i did is possible.. :p

OpenStudy (anonymous):

\[\sin(\frac{\pi}{3}) = ??\]

OpenStudy (anonymous):

\[\cos(\frac{\pi}{3}) = ??\]

OpenStudy (aroub):

Wait, the answer is 2?

OpenStudy (anonymous):

No.. Why can't you write it properly and check what is the answer you are getting? Be slow, are you running out of time?

OpenStudy (anonymous):

\[\frac{1}{\frac{\pi}{3}} \cdot \frac{\sin(\frac{\pi}{3})}{\cos(\frac{\pi}{3})}\]

OpenStudy (anonymous):

Be slow, evaluate this and tell me what did you get..

OpenStudy (anonymous):

Wait, I see, are you using this formula: \[\lim_{x \to 0} \frac{\sin(x)}{x} = 1\]

OpenStudy (aroub):

1.654

OpenStudy (aroub):

yesss, oh it doesn't apply here?

OpenStudy (anonymous):

Oh.. Be careful with the Limits given.. You are doing Limits, so you have to fully concentrate on what the limits are.. In this formula, x tending to 0 is must.. But in your original question, x is not tending to 0, are you getting this??

OpenStudy (anonymous):

Yes, 1.654 that you got is correct..

OpenStudy (aroub):

Oh yes! I see what i did wrong

OpenStudy (aroub):

Thank youuuuu :D

OpenStudy (anonymous):

You are welcome dear.. :)

OpenStudy (anonymous):

And for you knowledge: \[\lim_{x \to 0} \frac{\tan(x)}{x} = 1\]

OpenStudy (anonymous):

Again remember, in this x must tend to 0..

OpenStudy (aroub):

Haha, yes!

OpenStudy (anonymous):

Okay, that's great. :) Keep it up.. :)

OpenStudy (aroub):

I have another question, its a bit tricky. I'll post it now

OpenStudy (anonymous):

Sure..